OpenStudy (anonymous):

Let P = (a,b) lie in the first quadrant. Find the slope of the line through P such that the triangle bounded by this line and the axes in the first quadrant has minimal area. Then show that P is the midpoint of the hypotenuse of this triangle.

OpenStudy (amistre64):

use: y = m(x-a) + b area = xy/2 = x(m(x-a) + b)/2 area' = [ m(x-a) + b + mx ]/2 = 0 m(x-a) + b + mx = 0 and solve

OpenStudy (amistre64):

m(2x - a) + b = 0 m = -b/(2x-a)

OpenStudy (amistre64):

y = -b(x-a)/(2x-a) + b y = b [ 1 - (x-a)/(2x-a) ] y = b (2x - a - x +a)/(2x-a) y = bx/(2x-a) midpoint eh ....

OpenStudy (amistre64):

intercepts i spose would be needed